# thanks: Corresponding author

### v1

@@ Abstract
Introduction   Since the publication of general relativity (GR), Albert Einstein’s geometric theory of gravitation, in 1915, numerous alternative theories have been proposed to address its limitations or explore different formulations. These include the Brans-Dicke theory Brans and Dicke (1961); Dicke (1962), $f(R)$ gravity Buchdahl (1970); Starobinsky (1980); Sotiriou and Faraoni (2010); De Felice and Tsujikawa (2010), Modified Newtonian Dynamics (MOND) Milgrom (1983); Bekenstein (2004); Famaey and McGaugh (2012), massive gravity Fierz and Pauli (1939); de Rham et al. (2011); de Rham (2014), Horndeski theory Horndeski (1974); Gleyzes et al. (2015); Langlois (2019), Loop Quantum Gravity (LQG) Ashtekar (1986); Ashtekar and Lewandowski (2004); Perez (2013); Dapor and Liegener (2018), Causal Dynamical Triangulations (CDT) Ambjorn and Loll (1998); Ambjorn et al. (2012); Loll (2020); Ambjørn et al. (2019), and emergent gravity Verlinde (2011, 2017) (see Refs. Clifton et al. (2012); Joyce et al. (2015); Koyama (2016); Berti et al. (2015); Akrami et al. (2021) for reviews). In addition, Bañados and Ferreira proposed a modified theory of gravity known as Eddington-Inspired Born-Infeld (EiBI) gravity Banados and Ferreira (2010), which draws inspiration from the Born-Infeld electromagnetic action and Eddington’s purely affine formulation of gravity. This theory avoids cosmological singularities and modifies gravitational dynamics in high-density regimes, such as the interiors of compact astrophysical objects (e.g., black holes and neutron stars) and the early universe Pani et al. (2011, 2012); Harko et al. (2013); Avelino and Ferreira (2012); Avelino (2012a); Delsate and Steinhoff (2012); Scargill et al. (2012); Sham et al. (2013); Bouhmadi-López et al. (2015); Sotani and Miyamoto (2014); Wei et al. (2015); Du et al. (2014); Avelino (2016a, b); Albarran et al. (2019); Jiménez Cruz and Escamilla-Rivera (2019). It is worth noting in this context that EiBI gravity also predicts significant deviations from GR at subatomic scales, including within protons Avelino (2012b, 2019).
In EiBI theory of gravity, the effective gravitational pressure can lead to a significantly stronger gravitational field inside matter compared to GR Avelino (2012a); Delsate and Steinhoff (2012), despite both theories being identical in vacuum. In Ref. Avelino (2019), it has been demonstrated that constraints on EiBI’s additional parameter, $\kappa$, can be derived from the pressure distribution of quarks within the proton, as measured via deeply virtual Compton scattering (DVCS) data Burkert et al. (2018). The resulting bound, $|\kappa|\lesssim c^{4}|p_{\text{peak}}|^{-1}\sim 10^{-1}\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (where $c$ is the speed of light and $p_{\text{peak}}$ is the average peak pressure near the center of the proton), is nearly an order of magnitude stronger than limits inferred from neutron star core densities Pani et al. (2011) (and six order of magnitude stronger than limits inferred using solar physics Casanellas et al. (2012)), though weaker than those from other studies Avelino (2012a, b); Latorre et al. (2018); Delhom et al. (2020); Beltrán Jiménez et al. (2021). While the connection between the internal gravitational field strength and proton pressure distributions is interesting, further investigation is needed to assess how different theoretical or experimental determinations of the proton’s pressure profile might influence the constraints on $\kappa$.
@@
where $S_{M}(g,\Phi)$ represents the matter action and $\Phi$ refers to a general matter field. The parameter $\lambda$ is dimensionless and related to the cosmological constant via $\Lambda=(\lambda-1)/\kappa$, while $\kappa$ is an additional EiBI parameter with dimensions of length squared. Note that throughout this paper, we work in natural units where $18\pi G=c=$ 1, except when stated otherwise. EiBI gravity is formulated in the Palatini approach Olmo (2011), wherein the metric tensor and the affine connection are regarded as independent dynamical variables. The equations of motion can be derived by performing independent variations of action (1) with respect to the connection $\Gamma^{\rho}_{~{}\mu\nu}$ and the metric $g_{\mu\nu}$. This yields, respectively,
@@
In the context of EiBI gravity, the Poisson equation is modified to incorporate a correction term dependent on the matter distribution. The equation takes the form Pani et al. (2012):
$\nabla^{2}\phi(r)=4\pi G_{N}\rho(r)+\frac{\kappa}{4}\nabla^{2}\rho(r),$ (6)
where $\phi(r)$ is the gravitational potential, $\rho(r)$ is the energy density and $G_{N}$ is Newton’s gravitational constant. The additional term $\frac{\kappa}{4}\nabla^{2}\rho(r)$ introduces a correction to the Newtonian potential. Note that depending on the energy density distribution function $\rho(r)$, the behaviour of Laplacian acting on $\rho(r)$, and the sign of $\kappa$, the resulting interaction can manifest as either attractive or repulsive at different distances from the center.
@@
Conclusions   In this work, we have explored the constraints on the EiBI gravity parameter $\kappa$ using the proton’s internal pressure distribution. To this aim, we have used the pressure distributions determined by the MMGPDs Collaboration Goharipour et al. (2025a) through the QCD analysis of GPDs. Our study reveals that the bounds on $\kappa$ are sensitive to the specific proton pressure model employed. The stronger limits are obtained for higher values of parameter $M^{2}$ ($M^{2}=2\,\text{GeV}^{2}$), which governs the $t$-dependence of GFF $D(t)$, compared to lower ones ($M^{2}=1\,\text{GeV}^{2}$). The derived constraints, ranging from $|\kappa|\leq 0.10\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (for peak pressures) to $|\kappa|\leq 0.3\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (for averaged pressures), are competitive with existing bounds from neutron stars and astrophysical observations Pani et al. (2011).

### v2

@@ Abstract
[…]
In EiBI theory of gravity, the effective gravitational pressure can lead to a stronger gravitational field inside matter compared to GR Avelino (2012a); Delsate and Steinhoff (2012), despite both theories being identical in vacuum. To be more precise, EiBI represents a class of modified gravity theories that can predict moderate deviations from GR within the matter for appropriate values of the EiBI’s additional parameter, $\kappa$. The theory naturally reduces to GR in the limit $\kappa\rightarrow 0$. In Ref. Avelino (2019), it has been demonstrated that constraints on $\kappa$ can be derived from the pressure distribution of quarks within the proton, as measured via deeply virtual Compton scattering (DVCS) data Burkert et al. (2018). The resulting bound, $|\kappa|\lesssim c^{4}|p_{\text{peak}}|^{-1}\sim 10^{-1}\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (where $c$ is the speed of light and $p_{\text{peak}}$ is the average peak pressure near the center of the proton), is nearly an order of magnitude stronger than limits inferred from neutron star core densities Pani et al. (2011) (and six order of magnitude stronger than limits inferred using solar physics Casanellas et al. (2012)), though weaker than those from other studies Avelino (2012a, b); Latorre et al. (2018); Delhom et al. (2020); Beltrán Jiménez et al. (2021). While the connection between the internal gravitational field strength and proton pressure distributions is interesting, further investigation is needed to assess how different theoretical or experimental determinations of the proton’s pressure profile might influence the constraints on $\kappa$.
@@
where $S_{M}(g,\Phi)$ represents the matter action and $\Phi$ refers to a general matter field. The parameter $\lambda$ is dimensionless and related to the cosmological constant via $\Lambda=(\lambda-1)/\kappa$, while $\kappa$ is an additional EiBI parameter with dimensions of length squared. Note that throughout this paper, we work in natural units where $8\pi G=c=1$, except when stated otherwise. EiBI gravity is formulated in the Palatini approach Olmo (2011), wherein the metric tensor and the affine connection are regarded as independent dynamical variables. The equations of motion can be derived by performing independent variations of action (1) with respect to the connection $\Gamma^{\rho}_{~\mu\nu}$ and the metric $g_{\mu\nu}$. This yields, respectively,
@@
In the context of EiBI gravity, the Poisson equation is modified to incorporate a correction term dependent on the matter distribution. The equation takes the form Avelino (2019):
$\nabla^{2}\phi=\frac{\rho}{2}+\frac{\kappa}{4}\nabla^{2}\rho\,.$ (6)
where $\phi(r)$ is the gravitational potential and $\rho(r)$ is the energy density. The additional term $\frac{\kappa}{4}\nabla^{2}\rho(r)$ introduces a correction to the Newtonian potential. Note that depending on the energy density distribution function $\rho(r)$, the behaviour of Laplacian acting on $\rho(r)$, and the sign of $\kappa$, the resulting interaction can manifest as either attractive or repulsive at different distances from the center.
@@
Conclusions   In this work, we have explored the constraints on the EiBI gravity parameter $\kappa$ using the proton’s internal pressure distribution. To this aim, we have used the pressure distributions determined by the MMGPDs Collaboration Goharipour et al. (2025a) through the QCD analysis of GPDs. Our study reveals that the bounds on $\kappa$ are sensitive to the specific proton pressure model employed. The stronger limits are obtained for higher values of parameter $M^{2}$ ($M^{2}=2\,\text{GeV}^{2}$), which governs the $t$-dependence of GFF $D(t)$, compared to lower ones ($M^{2}=1\,\text{GeV}^{2}$). The derived constraints, ranging from $|\kappa|\leq 0.10\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (for peak pressures) to $|\kappa|\leq 0.3\,\text{m}^{5}\,\text{kg}^{-1}\,\text{s}^{-2}$ (for averaged pressures), are competitive with existing bounds from neutron stars and astrophysical observations Pani et al. (2011), though they remain several orders of magnitude weaker than those obtained from collider experiments Beltran Jimenez et al. (2018); Latorre et al. (2018).